OpenStudy (anonymous):

Question 12 Simplify open parentheses x to the 2 third power close parentheses to the 4 fifths power x to the 2 fifths power x to the 2 fifteenths power x to the 8 fifteenths power x to the 22 fifteenths power Question 13 Simplify i31 1 −1 −i i Question 14 Identify the degree of the polynomial x2y + 5xy3 - 7y2 + 2. 3 4 5

OpenStudy (whpalmer4):

Is that \[(x^{2/3})^{4/5}\]? \[(x^a)^b = x^{a*b}\]

OpenStudy (whpalmer4):

\[i^1 = i\]\[i^2 = -1\]\[i^3 = i^2*i = -1*i = -i\]\[i^4 = i^2*i^2 = -1*-1 = 1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so whats #12?

OpenStudy (whpalmer4):

I gave you the rule to figure it out.

OpenStudy (anonymous):

@math&ing001 i need help. you explain better...

OpenStudy (whpalmer4):

For #14, the degree of the polynomial is the highest degree of any of the terms. The degree of a term is the sum of the powers. Degree of \(x^5y^3 = 5+3=8\)

OpenStudy (whpalmer4):

Compare \[(x^{2/3})^{4/5}\]with \[(x^a)^b = x^{a*b}\] Match up the pieces. Follow the directions.

OpenStudy (math&ing001):

Consider a=2/3 and b=4/5.

OpenStudy (anonymous):

would it be x^8/15?

OpenStudy (math&ing001):

yeah !

OpenStudy (anonymous):

so whats i^31?

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

i^31=i^30 * i i^30=(i^2)^15 and you know that i^2=-1

OpenStudy (anonymous):

right....

OpenStudy (anonymous):

-1^15 ?

OpenStudy (math&ing001):

(-1)^15 = -1

OpenStudy (anonymous):

#14?

OpenStudy (math&ing001):

Find the term with the highest power. That's your degree

OpenStudy (whpalmer4):

@math&ing001 not necessarily true. The degree of a term is the sum of the exponents of all the variables in the term. Degree of \(a^3b^2=5\), so degree of \[c^4+a^3b^2\]is 5, not 4 as your procedure would indicate.